Please help me with conversion.thanks in advance. 4.5m/s = ______ km/h; 3.5g/cm^3 = _________ kg/m^3; Please include a little explanation.thanks so much
Answer by stanbon(75887)

4.5m/s = ______ km/h
----
4.5 (meters/second) * (1 km/1000 meters)(3600 seconds/1 hr)
----
= (4.5*1*3600)/(1*1000*1)(km/hr)
----
= 162 (km/hr)
--------------------------
3.5g/cm^3 = _________ kg/m^3
--------
3.5(grams/centimeters)^3 * (1 kilogram/1000 grams)^3
--------
= (3.5*1)/(10^9) (kg/m^3
--------
= 0.000000003.5 (kg/m^3)
============
Cheers,
Stan H.
----------------